OpenStudy (anonymous):

A bag contains 6 cherries, 3 oranges and 2 lemon candies you reach in and take 3 pieces of candy at random. the probability that you have 1cherry and 2 lemon candies is?

OpenStudy (kropot72):

P(cherry) = 6/11 P(1st lemon) = 2/10 P(2nd lemon) = 1/9 P(required combination) = \[\frac{6\times 2\times 1}{11\times 10\times 9}=you\ can\ calculate\]